OpenStudy (anonymous):

what are the vertex,focus ,directrix of 5x=(y-3)^2

OpenStudy (zpupster):

for vertex : we ask when is x zero when y is 3? so what is (x,y)??

OpenStudy (zpupster):

you can write in a standard form maybe the x , y zero's are easier to see 5(x-0)^2 = (y-3)^2 focus is y^2 = 4ax but ours is y^2 =5x so 4ax= 5x solve for a and the y will be the same (a,y) the directix will be -a